OpenStudy (anonymous):

Find f(-4) if f(x)= 2x^3+x^2-13 I Have: f(-4) 2x^3+x^2-13 2(-4)^3+-4^2-13 -128+-16-13 Now what?

OpenStudy (anonymous):

you add like terms :D so you do -128-16-13

OpenStudy (anonymous):

I got -125

OpenStudy (anonymous):

so -128 -16 -13 you add em?

OpenStudy (anonymous):

idk i got -157

OpenStudy (anonymous):

you could do (-128)+(-16)+(-13) if you add them

OpenStudy (anonymous):

I thought it was -128 + 16 - 13

OpenStudy (anonymous):

-128+-16-13 you thought the +-16 was just a plus 16 the negative cancels the positive out so it would just be negative

OpenStudy (anonymous):

got -157 again so is tht correct??

OpenStudy (anonymous):

I thought it (-4)^2 is 16 not -16. So then -128 +16 = -112 -13 = -125. But Blondie might be right lol

OpenStudy (anonymous):

allea is right! i didnt look at the prob right. (-4)^2 is positive 16 :D

OpenStudy (anonymous):

Yayy! lol

OpenStudy (anonymous):

So its only 157 not negative

OpenStudy (anonymous):

-125

OpenStudy (anonymous):

No its -125. Do u need me to explain it so you can understand how i got my answer?

OpenStudy (anonymous):

i figured it out tnx thou hun :)

OpenStudy (anonymous):

Lol ur welcome :p

OpenStudy (anonymous):

haha